There are 17 meters of molding around the base of a rectangular room 3.9 meters long. How wide is the room
Answer by mananth(16949)

molding = 17 meters
perimeter = 17 meters
2*(L+W) =17
/2
L+W=17/2
L+w= 8.5
...
Length = 3.9
plug value of length
3.9+W=8.5
-3.9
W=8.5-3.9
W=4.6 meters
..
CHECK
3.9+3.9+4.6+4.6 = 17
..
m.ananth@hotmail.ca